OpenStudy (anonymous):

Let g(x) be the reflection of f(x) = 3x + 2 in the x-axis. What is a function rule for g(x)? g(x) = 3x – 2 g(x) = –3x – 2 g(x) = 3x + 2 g(x) = –3x + 2

OpenStudy (helder_edwin):

if u have a point (x,y), the reflection in the x-axis transforms it into (x,-y)

OpenStudy (helder_edwin):

since u have y=f(x)=3x+2

OpenStudy (helder_edwin):

then (x,y)=(x,f(x))=(x,3x+2) transforms into \[ \large (x,-y)=(x,-f(x))=(x,g(x)) \]

OpenStudy (helder_edwin):

so g(x)=-f(x)

OpenStudy (anonymous):

so its either b or c?

OpenStudy (helder_edwin):

\[ \large g(x)=-f(x)=-(3x+2) \]